OpenStudy (anonymous):

-7(x-4) + 8x = 12 + 14 What is the value of 'x'?

OpenStudy (bradely):

-7(x-4) + 8x = 12 + 14 Using distributive property a(b-c)= ab- ac -7x+28+8x=26 x+28=26 x=-2 Source: http://www.mathskey.com/question2answer/